# AdaSteer: Your Aligned LLM is Inherently an Adaptive Jailbreak Defender

### v1

@@ D.3 Analysis on Adaptive Steering
Tables 7 and Table 8 display the $pos_{\text{RD}}$ and $pos_{\text{HD}}$ along with their respective $\lambda_{r}$ and $\lambda_{c}$, for each data type on Qwen2.5 and Gemma-2, respectively. On the RD, we consistently observe that more rejection vectors are effectively applied to input types with lower $pos_{\text{RD}}$. In contrast, on the HD, Qwen2.5 does not clearly differentiate the harmfulness of inputs compared to LLaMA-3.1 and Gemma-2, leading to similar $pos_{\text{HD}}$ for both jailbreak and benign inputs. However, due to further tuning on the validation set, AdaSteer still manages to perform well on Qwen2.5.
@@ D.4 Analysis on Steering Vector and Model Size
LLaMA-3.1 96.00 92.80 94.40
ROSE 91.30 89.60 90.45
Self-CD 94.70 92.80 93.75
Jailbreak Antidote 95.70 87.20 91.45
Surgical 90.30 74.40 82.35
InferAligner 85.30 75.60 80.45
CAST 96.00 94.00 95.00
AdaSteer (Ours) 97.30 98.40 97.85
Qwen2.5 94.00 96.00 95.00
ROSE 98.00 96.00 97.00
@@
Jailbreak Antidote 94.30 92.00 92.15
Surgical 93.70 96.80 95.25
InferAligner 94.00 92.80 93.40
CAST 96.00 95.20 95.60
AdaSteer (Ours) 87.00 95.20 91.10
Gemma-2 89.30 83.20 86.25
ROSE 80.70 82.80 81.75
Self-CD 87.70 82.80 85.25
Jailbreak Antidote 88.70 78.00 83.35
Surgical 90.10 90.80 90.45
InferAligner 83.70 65.20 74.45
CAST 80.70 83.20 81.95
AdaSteer (Ours) 92.00 93.60 92.80
@@
LLaMA-3.1 57 30 0 60 61 67 37 38.14 96.00 92.80 50.00
AdaSteer (Ours) 100 100 82 90 85 100 86 91.86 97.30 98.40 50.01
w/o $\boldsymbol{v}_{\text{RD}}$ 47 35 0 64 64 22 45 39.57 98.70 98.40 50.70
w/o $\boldsymbol{v}_{\text{HD}}$ 100 100 96 78 95 91 81 91.57 82.30 66.40 45.72
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 100 100 95 86 87 98 84 92.14 94.00 96.40 47.02
@@
w/o $\boldsymbol{v}_{\text{RD}}$ 25 73 23 90 46 14 51 46.00 94.70 98.40 47.82
w/o $\boldsymbol{v}_{\text{HD}}$ 100 100 76 96 92 100 86 92.86 76.00 83.20 36.37
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 100 100 58 100 83 100 71 87.43 88.70 92.40 48.05
Gemma-2 6 31 0 90 57 1 27 30.29 89.30 83.20 50.00
AdaSteer (Ours) 91 95 75 86 86 86 82 85.56 93.60 92.00 48.28
w/o $\boldsymbol{v}_{\text{RD}}$ 14 98 22 94 78 16 74 56.57 91.30 86.00 49.99
w/o $\boldsymbol{v}_{\text{HD}}$ 100 99 100 60 86 100 100 92.14 82.30 98.00 33.08
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 98 100 99 68 90 94 91 91.43 94.00 99.20 46.00
@@
Qwen2.5-3B 13 47 0 56 40 5 6 23.86 94.80 94.70 50.00
AdaSteer (Ours) 94 97 56 88 79 100 48 80.29 94.40 93.70 45.72

### v2

@@ D.3 Analysis on Adaptive Steering
Tables 8 and Table 9 display the $pos_{\text{RD}}$ and $pos_{\text{HD}}$ along with their respective $\lambda_{r}$ and $\lambda_{c}$, for each data type on Qwen2.5 and Gemma-2, respectively. On the RD, we consistently observe that more rejection vectors are effectively applied to input types with lower $pos_{\text{RD}}$. In contrast, on the HD, Qwen2.5 does not clearly differentiate the harmfulness of inputs compared to LLaMA-3.1 and Gemma-2, leading to similar $pos_{\text{HD}}$ for both jailbreak and benign inputs. However, due to tuning on the validation set, AdaSteer still manages to perform well on Qwen2.5.
Jailbreak Attack Over-Safety Utility
[…]
Gemma-2-27B 2 4 0 94 58 1 36 27.86 84.80 90.33 50.00
+ AdaSteer 100 100 86 98 80 97 87 92.57 89.20 95.33 47.29
Table 4: Evaluation of AdaSteer on the large-scale Gemma-2-27B-it across seven jailbreak attacks, two over-safety benchmarks, and a utility benchmark.
@@ D.4 Analysis on Steering Vector and Model Size
We further evaluate AdaSteer on Gemma-2-27B, one of the most recent and powerful open-weight LLMs. As shown in Table 4, the base model exhibits limited robustness under various jailbreak attacks, with an average Defense Success Rate (DSR) of only 27.86%. In contrast, AdaSteer dramatically boosts defense performance across all seven attack types, achieving a DSR of 92.57%.
Importantly, AdaSteer preserves model utility: it maintains high helpfulness on benign prompts (as measured by a 47.29% win rate on AlpacaEval) and avoids excessive refusals, with over-safety refusal rates (CR) on par with the baseline (e.g., 84.80% $\rightarrow$ 89.20% on XSTest and 90.33% $\rightarrow$ 95.33% on OKTest). These results confirm that AdaSteer generalizes well to larger-scale models, maintaining strong safety-performance trade-offs without requiring any additional fine-tuning.
D.5 Analysis of Multilingual Attacks
Multilingual attacks present complexity due to linguistic variability and diverse syntactic structures (Zhao et al., 2025). However, we observe that AdaSteer demonstrates significant improvements in this scenario across all evaluated models. Specifically, for multi-language jailbreak attacks, AdaSteer improves the defense success rate on: LLaMA-3.1, from 67% to 100%, Qwen-2.5, from 14% to 90% andGemma-2, from 1% to 86%. These results demonstrate AdaSteer’s strong adaptability and generalization in handling multilingual adversarial prompts. While we acknowledge there is still room for further enhancement, especially in low-resource language settings, the current results show that AdaSteer already provides a substantial boost in defense effectiveness compared to baselines.
Appendix E Further Discussion
E.1 Nonlinear Steering Mechanisms
Currently, AdaSteer is built upon the widely adopted linear representation theory of activation space in LLMs (Zou et al., 2023a; Park et al., 2024), which assumes that certain behavioral features (e.g., harmfulness or rejection) can be captured through linear directions. While nonlinear steering mechanisms may further enhance control and expressivity, their theoretical foundations and practical implementations remain largely unexplored and unvalidated in the context of activation-based researches.
E.2 Combined with Training-related Strategies
We believe that AdaSteer can indeed be effectively combined with training-based strategies to further enhance both security and utility. One promising direction would be to treat the AdaSteer-modified representations at each layer as target labels, and the original model’s representations as inputs, using a mean squared error (MSE) loss to fine-tune the model directly toward the desired behavior.
This would allow the model to internalize AdaSteer’s behavior as part of its own parameters, potentially reducing inference-time overhead while preserving its defensive effectiveness.
E.3 Limited Probing Data
Regarding the number of compiled harmful examples, we are surprised to find that even a small number of such examples is sufficient to achieve the desired effect. In our main experiments, we use 15 compiled harmful examples for LLaMA-3.1-8B-Instruct, 13 for Qwen2.5-7B-Instruct, and 5 for Gemma-2-9B-it. In addition, we include an equal number of rejected harmful examples and complied benign data for each model. In our experiments, we found that even with such limited data, AdaSteer is able to identify meaningful harmful directions and achieve strong defense performance across a range of jailbreak attacks. This demonstrates the method’s data efficiency and practicality, especially in scenarios where access to large-scale harmful data is limited.
E.4 On the Plug-and-Play Property of AdaSteer
Once the Rejection Direction (RD) and Harmfulness Direction (HD) are extracted, we do not perform any additional adjustments for different attack types or data distributions. One of the core strengths of AdaSteer is that these directions, once computed, remain fixed and reusable across diverse scenarios. As shown in Table 1, AdaSteer demonstrates strong robustness against a wide range of jailbreak strategies—including prompt injection, role-play attacks, and multilingual attacks—without the need to modify RD or HD. This validates the general applicability of the extracted directions and supports our claim that AdaSteer can serve as a plug-and-play defense mechanism across different threat models.
@@ E.4 On the Plug-and-Play Property of AdaSteer
LLaMA-3.1 92.80 96.00 94.40
ROSE 89.60 91.33 90.47
Self-CD 92.80 94.67 93.74
Jailbreak Antidote 87.20 95.67 91.44
Surgical 74.40 90.33 82.37
InferAligner 75.60 85.33 80.47
CAST 94.00 96.00 95.00
AdaSteer (Ours) 98.40 97.33 97.87
Qwen2.5 96.00 94.00 95.00
ROSE 96.00 98.00 97.00
@@
Jailbreak Antidote 92.00 94.33 93.17
Surgical 96.80 93.67 95.24
InferAligner 92.80 94.00 93.40
CAST 95.20 96.00 95.60
AdaSteer (Ours) 95.20 87.00 91.10
Gemma-2 83.20 89.33 86.27
ROSE 82.80 80.67 81.74
Self-CD 82.80 87.67 85.24
Jailbreak Antidote 78.00 88.67 83.34
Surgical 90.80 90.33 90.57
InferAligner 65.20 83.67 74.44
CAST 83.20 80.67 81.94
AdaSteer (Ours) 93.60 92.00 92.80
@@
LLaMA-3.1 57 30 0 60 61 22 37 38.14 92.80 96.00 50.00
AdaSteer (Ours) 100 100 82 90 85 100 86 91.86 98.40 97.33 50.01
w/o $\boldsymbol{v}_{\text{RD}}$ 47 35 0 64 64 22 45 39.57 98.40 98.67 50.70
w/o $\boldsymbol{v}_{\text{HD}}$ 100 100 96 78 95 91 81 91.57 66.40 82.33 45.72
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 100 100 95 86 87 98 84 92.14 96.40 94.00 47.02
@@
w/o $\boldsymbol{v}_{\text{RD}}$ 25 73 23 90 46 14 51 46.00 98.40 94.67 47.82
w/o $\boldsymbol{v}_{\text{HD}}$ 100 100 76 96 92 100 86 92.86 83.20 76.00 36.37
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 100 100 58 100 83 100 71 87.43 92.40 88.67 48.05
Gemma-2 6 31 0 90 57 1 27 30.29 83.20 89.33 50.00
AdaSteer (Ours) 91 95 75 86 86 86 82 85.56 92.00 93.67 48.28
w/o $\boldsymbol{v}_{\text{RD}}$ 14 98 22 94 78 16 74 56.57 86.00 91.33 49.99
w/o $\boldsymbol{v}_{\text{HD}}$ 100 99 100 60 86 100 100 92.14 98.00 82.33 33.08
w/ reverse $\boldsymbol{v}_{\text{RD}}$ 98 100 99 68 90 94 91 91.43 99.20 94.00 46.00
@@
Qwen2.5-3B 13 47 0 56 40 5 6 23.86 94.80 94.67 50.00
AdaSteer (Ours) 94 97 56 88 79 100 48 80.29 94.40 93.67 45.72